OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2.

OpenStudy (goformit100):



OpenStudy (whpalmer4):

@jhannybean @goformit100 What does a directrix of y=-2 mean to you?

OpenStudy (whpalmer4):

@jhannybean @goformit100 Perhaps a sketch showing both your proposed parabola and y = -2 would be informative :-)

OpenStudy (whpalmer4):

Just draw me a picture, okay? :-)

OpenStudy (whpalmer4):

Draw me a picture showing y = -2 as the directrix with the parabola you specified...

OpenStudy (whpalmer4):

Okay, I'll plot it.

OpenStudy (whpalmer4):

I mislabeled the parabola — though it says \(y^2=x\) it is a graph of \(y^2=8x\)

OpenStudy (jhannybean):

What do you use to show your graphs btw?

OpenStudy (whpalmer4):

Mathematica

OpenStudy (jhannybean):

Ahh, ok.

OpenStudy (whpalmer4):

Does \(y=-2\) look like the directrix of that parabola? :-)

OpenStudy (jhannybean):

Oh i see...it would have been x=-2.... Hmm.

OpenStudy (jhannybean):

I'll delete my work. lol.

OpenStudy (whpalmer4):

Why not post a corrected version?

OpenStudy (jhannybean):

http://www.sketchtoy.com/39052824 he distance from the focus to the vertex is equal to the distance from the directrix to the vertex. The distance is labeled as "p" Use the form \[\large (x-h)^2 = 4p(y-k)\] since this parabola is opening up and down. vertex= (0,0) ,p = 2 (distance of center from focus and center to directrix) \[\large (x-0)^2 = 4(2)(y-0)\]\[\large x^2 = 8y\] And your final graph will look something like http://www.sketchtoy.com/39073173